Atlas Sw Validation Reports

Reports of the Software Validation on Tuesday April 24, 2007

Manuel Gallas

Wednesday April 24, 2007

Phone Meeting

The next phone meeting will be on Tuesday, May 15, 2007 at 16:10 CERN time (from 16:10 to 17:40). The 1st May is holiday day at CERN (and in most of the countries) the 8th of May is again the time slot for the physics validation meeting.

The hour of the meeting will be probably changed to 16:30 pm to allow people to attend SIT and sw-validation meetings. It will be announced by e-mail if this is the case.

Meeting Coordinates:

ATLAS software validation

(Manuel Gallas)

Dial-in numbers: +41227676000 (Main) Access codes: 0132753 (Participant) Participant site: https://audioconf1.cern.ch/call/0132753

IMPORTANT: During this 3 weeks period without sw-validation phone meeting please send the report of your domain/detector in a written format preferably before Monday at noon (in this way the report to the SPMB will contain the most updated information).


Software validation coordination

Report on the actions from the previous meeting:

Actions:*

  • Create a hyper news forum for the physics and software validation activities. (DONE, https://hypernews.cern.ch/HyperNews/Atlas/get/physics-software-validation.html)
  • Create a mailing list for the communication with the software validation coordinators (DONE, all the sw-validation coordinators included).
  • Create a Twiki page to host the links of the software validation activities in the different software domains and sub-detector software. (DONE, initial skeleton linked from the AtlasReleasValidation https://twiki.cern.ch/twiki/bin/view/Atlas/SoftwareValidation)
  • Follow up those items described in the A.O.B section: RTT and ATN documentation and examples, the mailTo , doc, and rtt classification tags, counting the ERROR lines using the existing RTT machinery.

(ongoing)


Report from the physics validation activities


Report from the software domain and sub-detector software

Core Services (Paolo Calafiura)

Status:

We are now using (apparently successfully) LCG51 and the matching Gaudi v19r2p3.

News:

A new v19r2p4 with a fix from Sebastien to speedup athena startup (by "hiding" the individual packages rootmap files from ROOT). This will go in 13.x.0 and once tested will be moved to 13.0.x.

Database (David Malon)

No report.

Infrastructure (Fred Luehring)

Right now there are no open Software infrastructure issues for release 13

Generators (Giorgos Stavropoulos)

No report

Simulation (Adele Rimoldi)

Status:

the atlas core simulation seems OK in the ATN tests but the RTT tests are not running as good as expected: sometimes there is a timeOut by the system, in some jobs we had problems with castor and in some other cases the test configuration seems corrupted (investigations ongoing).

Digitization (Sven Vahsen)

Issues:

  • RDO percistency is still not fully there. (This relevant packages are in AtlasEvent, not in AtlasSimulation.) Ed Moyse has a fix, but seems hesitent to put it in the release, as he's in the middle of some other changes in one of the packages that contains the fix. (see http://savannah.cern.ch/bugs/?25670).
  • LVL1 jobOptions in digitization had become outdated, but trigger people have been working on it. Hope this is OK in next nightly.
  • Ranlux random service shows strange behavior, we keep old ranecu based service as default for now.

Reconstruction (David Rousseau)

No report

EDM (Davide Costanzo, RD Schaffer)

News:

from the validation perspective there is the request to be able to read release 13 hits with release 12 digitization to validate release 13 simulation (and the new g4 version).

PAT (Ketevi Assamagan, Tadashi Maeno)

No report

Inner Detector (Markus Elsing)

No report

LAr Calorimeter (Hong Ma, Guillaume Unal, Karim Bernadet (TBC))

Status:

  • Looking at the RTT results for the Calo only in rel_6, bugfix and i686-slc4-gcc34-opt simulation and digitization look ok reco ( and with ESD and AOD, ESD then AOD ~works too) looks ok.
  • Concerning the DetFlags.LVL1_setOn() problem I had switch off the LVL1 in the RTT jobs, but it should be fixed https://savannah.cern.ch/bugs/?25584. I didnt check it (will try today)
  • There are errors in reading an ESD to do AOD.
  • In the reco step (if you looks at the CaloRecEx results), initialization times still take a long time

Tile Calorimeter (Sasha Solodkov)

Status

is essentially the same as LAr, because Karim is looking at combined Calo tests. Sasha will monitor the tests from TileCal side to resolve problems which are TileCal-specific.

Muon Spectrometer (Steve Goldfarb, Lashkar Kashif)

Status:

ATN

AtlasSimulation: All packages now build correctly.

AtlasReconstruction: The following four packages don't build: CSC_DHoughSegmentMakerAlg, CSC_DHoughSegmentMakerTool (the developer expects to solve the problems by Wednesday April 25), MuidParticleCreator (waiting for fix in IIsolationTool by Ketevi. New tag for IsolationTool has been submitted), MuonTB Ana (Haven't heard back from package manager).

AtlasAnalysis: One package fails to build:

MuonMonitoring

Haven't heard back from package manager.

The muon packages are still failing their RTTs. I heard back from Stephane concerning MooPerformance:

"We are still recovering from the EDM changes, configurables, etc. The software is still broken for the full chain. I need to set up more restricted tests also, e.g. just running Moore inside the Muon Spectrometer and avoid ESD/AOD making." Nectarios is also looking into this issue.

Tracking (Wolfgang Liebig)

No report.

Trigger ( Ricardo Goncalo, Olga Igonkina)

Status:

Electrons

  • only e10 works: this is only through missing configuration (migration to configurables) and should be fixed in tonight's nightly
  • it works up to EF except for the hypo after TrigEgammaRec
  • L2: fully configurable
  • EF: TrigCaloRec maybe not in configurables yet
  • EF: InDet fully configurable now
  • EF: TrigEgammaRec not migrated yet
  • monitoring histograms: only in L2CaloHypo and IdScan

Photons:

  • needs to migrate to configurables: being done this week
  • implementing monitoring histos this week

Muons:

  • Status: not working (this has probably improved by now)
  • new tags from Alessandro should fix the current problems
  • monitoring histos: planning to have this asap but very busy at the moment

Tau:

  • L2: runs
  • EF: changes in EDM affecting taus but hope to have it working soon
  • configurables in L2 and almost done in EF
  • monitoring histos implemented in HLT

B-tag:

  • L2: b45 runs in recent nightles
  • migrated to configurables
  • first monitoring histo added, more to follow

Cosmics:

  • L2 compiling fine in nightlies
  • migrating to new monitoring histograms

L2 InDet:

  • new monitoring histograms implemented in IdScan, SiTrack should follow

L2 Calo:

  • code running in nightlies
  • has old histograms, not yet new framework

Jets:

  • 12.0.6 looks ok
  • configurables working on nightlies for L2 and EF
  • TrigCaloRec is more complicated (involves several slices)
  • monitoring histos: some histos in L2 and EF jets (working in 13.0.x)
  • ideas: variables to check being identified for monitoring and validation

Bphys:

  • not running in recent nightlies
  • problem from un-seeded algos: being investigated
  • No monitoring histos yet in repository (but work ongoing)

EF InDet:

  • Monitoring histos: using old monitoring packages, no monitoring histos added yet
  • ATN test: 05IDtransBS active in nightlies for IdScan, SiTrack and EF InDet

RTT tests:

TrigEgammaValidation: some tests inactive now (depends on atlasAnalysis package which is broken) one test is working TrigInDetRecExample (new) saves CBNT plus histograms

L1 Muons:

  • Muon group produces AANT file for studies: can be used for monitoring

L1 CTP:

  • muCTPI & CTP simulations are migrated to configurables
  • bytestream monitoring package may be usable for dataset validation

Other:

  • TrigTimerSvc not working: fix very soon

Production Transforms (Manuel Gallas)

Status:

  • The 12.0.6.4 cache is in production. One of the internal tests included with the cache fails when we try to produce TAG, SAN, MergedAOD, and HPTV at the same time. This failure is not reproduced by the Full Chain Test in which we use different input files.
  • Another issue is that recoESD seems not working if Truth is Off (investigation ongoing)
  • The 12.0.6.5 nightlies open, deadline for the new cache 2nd May
  • Activation of the 13.0.0 production scripts starting right now.

A.O.B


-- Main.gallasm - 25 April 2007

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-05-06 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback